# How weak are weak factors? Uniform inference for signal strength in signal plus noise models

### v1

@@ 9. Appendix B: Applications of general theorem to four models
In this section we use Theorem 8.13 to prove Theorem 4.6. The proofs for all four settings follow similar outlines, and we present the most detailed exposition for spiked Wigner matrices. The difficulty of the argument increases steadily from the Wigner to the CCA case, and we recommend sequential reading.
@@ 9.1. Spiked Wigner matrices
For the spiked Wigner case of Theorem 4.6, our first task is to introduce an equation governing the change of the eigenvalues of a symmetric matrix under a rank one perturbation.
Take $N\times N$ real symmetric matrix $\mathbf{B}$ with eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N}$ and normalized eigenvectors $\mathbf{u}_{i}$, $1\leq i\leq N$, satisfying $\langle\mathbf{u}_{i},\mathbf{u}_{j}\rangle=\delta_{i=j}$. In addition, take a column-vector $\mathbf{u}^{*}$. For a real constant $\theta\neq 0$ we define
@@
Assuming without loss of generality that all $\lambda_{i}$ are distinct and for all $i$ we have $\langle\mathbf{u}_{i},\mathbf{u}^{*}\rangle\neq 0$ (general case is obtained by a limit transition), note that (9.1) is a polynomial equation on $a$ of degree $N$, and therefore has $N$ roots which are $\mu_{1}$, …, $\mu_{N}$. Tracking the sign changes of the difference between the RHS and the LHS on intervals $(-\infty,\lambda_{N})$, $(\lambda_{N},\lambda_{N-1})$,…, $(\lambda_{1},+\infty)$, we localize the roots, so that they satisfy the desired interlacements. ∎
The plan for the rest of the section is to recursively use (9.1) to produce an inductive proof of Theorem 4.6 for the spiked Wigner model by growing $r$. On each step, we need to know the law of the scalar products $\langle\mathbf{u}_{i},\mathbf{u}^{*}\rangle$ appearing in the equation. This is simplified by the following observation:
@@
One can replace the vectors $\mathbf{u}_{i}^{*}$ in (2.1) with $r$ vectors uniformly distributed on the $N$–dimensional unit sphere, independent from each other and from $\mathcal{E}$: Theorem 4.6 for deterministic orthonormal vectors as in Section 2.1 is equivalent to the same theorem for independent unit vectors.
@@
We first note that the asymptotics in Theorem 4.6 does not depend on the choice of the unit vectors $\mathbf{u}_{i}^{*}$ in (2.1) as long as they are orthogonal to each other. Indeed, any $r$ orthogonal vectors can be obtained from any other $r$ orthogonal vectors by an orthogonal transformation of the space. Such transformation does not change the eigenvalues of $\mathbf{A}$, and also does not change the probability distribution of the matrix $\mathcal{E}$ (which uses Gaussianity of its matrix elements).
@@
(9.5) $\lim_{N\to\infty}\sqrt{N}(\lambda_{i}-\lambda(\theta_{i}))$ =$\stackrel{{\scriptstyle d}}{{=}}\mathcal{N}(0,V(\theta_{i})),\qquad 1\leq i\leq q-1,$
(9.6) $\lim_{N\to\infty}N^{2/3}(\lambda_{i}-2)$ =$\stackrel{{\scriptstyle d}}{{=}}\mathfrak{a}_{i-q+1},\qquad i\geq q,$
where $\mathcal{N}(0,V(\theta_{i}))$, $10\leq i\leq q-$ 1, are independent and $\{\mathfrak{a}_{j}\}_{j\geq 1}$ are points of the Airy1 point process independent from (9.5). In addition, (8.46) with $h(x)=1$ holds for the eigenvalues of $\mathbf{A}$.
@@
[…]
@@
We proceed by induction on $r$; the base case $r=0$ is Proposition Proposition. We analyze the equation (9.1) with $\theta=\theta_{1}$ and $\lambda_{i}$ being eigenvalues of $\mathbf{A}=\sum_{i=2}^{r}\theta_{i}\cdot\mathbf{u}^{*}_{i}(\mathbf{u}^{*}_{i})^{\mathsf{T}}+\mathcal{E}$. We start by considering the case $\theta_{1}>\theta_{c}$ (i.e., $q>1$) and comment on the changes for the case $q=1$ at the end. We first look at the interval $a\in[\lambda_{1},+\infty)$. The right-hand side of the equation (9.1) is a monotone-decreasing function of $a$ in this interval, changing from +$\infty$+ ∞ to 00. Hence, there is a unique $\hat{a}\in(\lambda_{1},+\infty)$ solving (9.1). In order to locate this $\hat{a}$, we make an asymptotic expansion of the equation. Using Lemma 9.4, we can assume $\mathbf{u}^{*}$ in (9.1) to be a uniformly random vector on the unit sphere, independent from everything else. Then $\langle\mathbf{u}_{i},\mathbf{u}^{*}\rangle$ are coordinates of a similar vector, because $\mathbf{u}_{i}$ are orthonormal. Hence, the equation (9.1) is recast as
@@
For the second term, note that $\zeta_{j}^{2}-\tfrac{1}{N}$, $j=1,2,\dots,N$, are weakly dependent mean 00 random variables. Hence, conditionally on $\lambda_{1},\dots,\lambda_{N}$, CLT applies and the sum is asymptotically Gaussian. The limit of the variance can be computed using
@@
For the remaining eigenvalues, the idea is to show that the $(i+1)$st largest root of (9.8) is very close to $\lambda_{i}$ and then use the induction hypothesis. By Corollary 9.3, the $(i+1)$st largest root is the unique root in the interval $(\lambda_{i+1},\lambda_{i})$ and our task is to show that it is much closer to the right end-point of this segment rather than to the left end-point.
@@
Note that for $a$ close to $\lambda_{i}$, the value of $\frac{1}{2}\bigl{(}a-\sqrt{a^{2}-4}\bigr{)}$ is close to $1/\theta_{i+1}>1/\theta_{1}$ by induction assumption (9.5). Hence, for large $N$ the left-hand side of (9.13) is negative and bounded away from zero. On the other hand $\zeta_{i}^{2}=O(\tfrac{1}{N})$. We conclude that $a$ should be smaller than $\lambda_{i}$, at distance $O(\tfrac{1}{N})$, in order for (9.13) to hold. Hence, this root $a$ satisfies the asymptotics (9.5).
Case 2: $\theta_{i+1}<\theta^{c}$, so that $\lambda_{i}$ is close to $\lambda_{+}$. We approximate the right-hand side of (9.8) as $N\to\infty$ near $\lambda_{+}$ by writing $\zeta_{j}^{2}\stackrel{{\scriptstyle d}}{{=}}\frac{\xi_{j}^{2}}{\sum_{\ell=1}^{N}\xi_{\ell}^{2}}$ with i.i.d. $\mathcal{N}(0,1)$ random variables $\xi_{j}$ and then using Theorem 8.20 with $\mathfrak{s}=\mathfrak{m}=1$, $h(x)=1$, whose assumptions hold by the induction hypothesis. The right-hand side of (9.8) has asymptotics $1+N^{-1/3}\cdot\mathcal{G}(b)+o(N^{-1/3})$ where the rescaled variable is $b=N^{2/3}(a-2)$. The interval $a\in(\lambda_{i+1},\lambda_{i})$ turns asymptotically into $b\in(\mathfrak{a}_{i-q+2},\mathfrak{a}_{i-q+1})$, and the equation becomes $\mathcal{G}(b)=N^{1/3}(1/\theta_{1}-1)+o(N^{-1/3})$. Since $11/\theta_{1}<1/\theta^{c}=$ 1, we are looking for a value of $b\in(\mathfrak{a}_{i-q+2},\mathfrak{a}_{i-q+1})$, where $\mathcal{G}(b)$ would be large and negative. Clearly, then $b$ needs to be close to the right end-point of the segment, i.e., to $\mathfrak{a}_{i-q+1}$ and we achieve (9.6).
@@
It remains to compute the asymptotics of $a_{q}$. Because $\mathbf{u}^{*}$ is uniformly random and $[\mathbf{u}_{i}]_{i=1}^{N}$ are orthonormal, the vector $\langle\mathbf{u}_{j},\mathbf{u}^{*}\rangle^{2}$, $1\leq j\leq N$, has the same distribution as $\frac{\xi_{j}^{2}}{\sum_{\ell=1}^{N}\xi_{\ell}^{2}}$ with i.i.d. $\mathcal{N}(0,1)$ random variables $\xi_{j}$. Hence, recalling $\theta_{q}=\theta^{c}+N^{-1/3}\tilde{\theta}=1+N^{-1/3}\tilde{\theta}$, (9.1) becomes:
@@
Taylor expanding the left-hand side in small $N^{-1/3}\tilde{\theta}$ and using Corollary 8.22, we conclude that the desired root $b$ converges towards the largest root of the equation $-\tilde{\theta}=\mathcal{G}(b)$. Comparing with Definition 4.2, we are done. ∎
@@ 9.2. Spiked covariance model
The proof of Theorem 4.6 for the spiked covariance model of Section 2.2 follows the same outline as the argument in the previous section, but with a different computational part. We explain the new computations, but otherwise try being brief on the technical details which repeat the previous section. We start with an analogue of Proposition 9.1.
Suppose that we are given $N\times S$ matrix $\mathbf{U}$, in which the rows are indexed by $i=0,1,\dots,N-1$ and $S\geq N$. We let $\lambda_{*}$ denote the squared length of the zeroth row of $\mathbf{U}$ (treated as an $S$–dimensional vector) and let $\mathbf{v}^{*}$ be the unit vector in the direction of this row. Let $\widetilde{\mathbf{U}}$ denote the $(N-1)\times S$ matrix formed by rows $i=1,2,\dots,N-1$ of $\mathbf{U}$.
We would like to connect the singular values and singular vectors of $\mathbf{U}$ to: singular values and vectors of $\widetilde{\mathbf{U}}$,  $\lambda_{*}$,  and $\mathbf{v}^{*}$. We let $(\mathbf{u}_{i},\mathbf{v}_{i},\sqrt{\lambda_{i}})$, $11\leq i\leq N-$ 1, be the left singular vector (of $(N-1)\times 1$ dimensions), right singular vector (of $S\times 1$ dimensions), singular value triplets for $\widetilde{\mathbf{U}}$, which means that
@@
In order to see how Proposition 9.6 is relevant for the setting of Section 2.2, we choose one index $1\leq k\leq r$ and a deterministic orthogonal matrix $O$, which maps $\mathbf{u}_{k}^{*}$ into the zeroth basis vector and orthogonal complement of $\mathbf{u}_{k}^{*}$ to the span of the basis vectors with labels $11,2,\dots,N-$ 1, so that
@@
where $\Omega^{\prime}$ is $(N-1)\times(N-1)$ symmetric matrix with eigenvalues $\{\theta_{i}\}_{i\neq k}$ and $1$ of multiplicity $N-r$. Conjugating with $O$ does not change the eigenvalues of the sample covariance matrix $\frac{1}{S}XX^{\mathsf{T}}$; it also does not change the fact that the columns of $X$ are i.i.d. On the other hand, after transformation by $O$, we can use Proposition 9.6 for $\mathbf{U}=\Omega X$. We reach the following conclusion:
@@
On the other hand, the value of $\lambda^{*}$ in (9.18) is the squared length of the zeroth row, $\|\mathbf{v}^{*}\|^{2}=\sum_{l=1}^{S}\theta_{k}\xi_{l}^{2}$. Finally, Proposition 9.6 deals with $\mathbf{U}\mathbf{U}^{\mathsf{T}}$, while in the corollary the matrix $XX^{\mathsf{T}}$ is divided by $S$ and, therefore, we should rescale by $S$ both $\lambda_{i}$ and $a$. Hence, dividing (9.18) by $a$ and $\theta_{k}$, we get
$\frac{1}{a}\left(\frac{\sum_{i=1}^{S}\xi_{i}^{2}}{S}+\frac{1}{S}\sum_{i=1}^{N-1}\frac{\lambda_{i}\xi_{i}^{2}}{a-\lambda_{i}}\right)=\frac{1}{\theta_{k}}.$
Rearranging the terms, we arrive at (9.19). ∎
@@
(9.20) $\lim_{N\to\infty}\sqrt{N}(\lambda_{i}-\lambda(\theta_{i}))$ =$\stackrel{{\scriptstyle d}}{{=}}\mathcal{N}(0,V(\theta_{i})),\qquad 1\leq i\leq q-1,$
(9.21) $\lim_{N\to\infty}N^{2/3}\frac{\lambda_{i}-(1+\gamma)^{2}}{\gamma(1+\gamma)^{4/3}}$ =$\stackrel{{\scriptstyle d}}{{=}}\mathfrak{a}_{i-q+1},\qquad i\geq q,$
where $\mathcal{N}(0,V(\theta_{i}))$, $10\leq i\leq q-$ 1, are independent and $\{\mathfrak{a}_{j}\}_{j\geq 1}$ are points of the Airy1 point process independent from (9.20). In addition, (8.46) with $h(x)=1$ holds for the eigenvalues of the sample covariance matrix $\frac{1}{S}XX^{\mathsf{T}}$.
@@
The final statement, (8.46) with $h(x)=1$ follows by induction on $r$ from Theorem 8.24 for $r=0$, and using Lemma 8.28 with Corollary 9.8 for the induction step. Here $\lambda_{\pm}=(1\pm\gamma)^{2}$,
[…]
@@
We analyze the equation (9.19) for $k=q$. In this situation the asymptotics of $(\lambda_{i})_{i=1}^{N}$ is given to us by Proposition 9.9. Arguing as in the previous section, the $q-1$ largest roots of the equation are close to $\lambda_{1},\dots,\lambda_{q-1}$, resulting in (4.6). In order to establish (4.7), we need to approximate (9.19) for $a$ close to $\lambda_{+}=(1+\gamma)^{2}$ and locate the root of the equation in the $(\lambda_{q},\lambda_{q-1})$ interval. We change the variables
@@
Recognizing the constant $\kappa_{2}$ in the right-hand side of the last formula and comparing with Definition 4.2, we arrive at (4.7). ∎

### v2

@@ 9. Appendix B: Applications of general theorem to four models
In this section we use Theorem 8.13 to prove Theorem 4.6. The proofs for all four settings follow similar outlines, and we present the most detailed exposition for spiked Wigner matrices. The difficulty of the argument increases steadily from the Wigner to the CCA.
@@ 9.1. Spiked Wigner matrices
Our first task is to introduce an equation governing the change of the eigenvalues of a symmetric matrix under a rank one perturbation.
Take $N\times N$ real symmetric matrix $\mathbf{B}$ with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{N}$ and normalized eigenvectors $\mathbf{u}_{i}$, $1\leq i\leq N$, satisfying $\langle\mathbf{u}_{i},\mathbf{u}_{j}\rangle=\delta_{i=j}$. In addition, take a column-vector $\mathbf{u}^{*}$. For $\theta\neq 0$, set
@@
Assuming that all $\lambda_{i}$ are distinct and for all $i$ we have $\langle\mathbf{u}_{i},\mathbf{u}^{*}\rangle\neq 0$ (general case is obtained by a limit transition), note that (9.1) is a polynomial equation on $a$ of degree $N$, and therefore has $N$ roots which are $\mu_{1}$, …, $\mu_{N}$. Tracking the sign changes of the difference between the RHS and the LHS on intervals $(-\infty,\lambda_{N})$, $(\lambda_{N},\lambda_{N-1})$,…, $(\lambda_{1},+\infty)$, we localize the roots, so that they satisfy the desired interlacements. ∎
In the rest of the section we use (9.1) to produce an inductive proof of Theorem 4.6 for the spiked Wigner model by growing $r$. On each step, we need to know the law of the scalar products $\langle\mathbf{u}_{i},\mathbf{u}^{*}\rangle$ appearing in the equation. This is simplified by the following observation:
@@
In (2.1), replacing the deterministic orthonormal vectors $\mathbf{u}_{i}^{*}$ by $r$ independent vectors uniformly distributed on the NN–dimensional unit sphere (independent of $\mathcal{E}$) leaves Theorem 4.6 unchanged.
@@
We first note that the asymptotics in Theorem 4.6 does not depend on the choice of $\mathbf{u}_{i}^{*}$ in (2.1) as long as they are orthogonal to each other. Indeed, any $r$ orthogonal vectors can be obtained from any other $r$ orthogonal vectors by an orthogonal transformation of the space. Such transformation does not change the eigenvalues of $\mathbf{A}$, and also does not change the probability distribution of the matrix $\mathcal{E}$ (which uses Gaussianity of its matrix elements).
@@
(9.6) $\lim_{N\to\infty}N^{2/3}(\lambda_{i}-2)$ $\stackrel{{\scriptstyle d}}{{=}}\mathfrak{a}_{i-q+1},\qquad i\geq q,$
where $\mathcal{N}(0,V(\theta_{i}))$ are independent over $i$ and with points of the Airy1 point process $\{\mathfrak{a}_{j}\}_{j\geq 1}$. In addition, (8.45) with $h(x)=1$ holds for the eigenvalues of $\mathbf{A}$.
@@
(9.7) $\mu(x)\mathrm{d}x=\frac{1}{2\pi}\sqrt{4-x^{2}}\,\mathbf{1}_{[-2,2]}\,\mathrm{d}x,\qquad m(z)=\frac{z-\sqrt{z^{2}-4}}{2},$
@@
We proceed by induction on $r$; the base case $r=0$ is Proposition Proposition. We analyze the equation (9.1) with $\theta=\theta_{1}$ and $\lambda_{i}$ being eigenvalues of $\mathbf{A}=\sum_{i=2}^{r}\theta_{i}\cdot\mathbf{u}^{*}_{i}(\mathbf{u}^{*}_{i})^{\mathsf{T}}+\mathcal{E}$. We start by considering the case $\theta_{1}>\theta_{c}$ (i.e., $q>1$) and comment on the changes for the case $q=1$ at the end. We first look at the interval $a\in[\lambda_{1},+\infty)$. The right-hand side of the equation (9.1) is a monotone-decreasing function of $a$ in this interval, changing from $\infty$ to 0. Hence, there is a unique $\hat{a}\in(\lambda_{1},+\infty)$ solving (9.1). In order to locate this $\hat{a}$, we make an asymptotic expansion of the equation. Using Lemma 9.4, we can assume $\mathbf{u}^{*}$ in (9.1) to be a uniformly random vector on the unit sphere, independent from everything else. Then $\langle\mathbf{u}_{i},\mathbf{u}^{*}\rangle$ are coordinates of a similar vector, because $\mathbf{u}_{i}$ are orthonormal. Hence, the equation (9.1) is recast as
@@
For the second term, note that $\zeta_{j}^{2}-\tfrac{1}{N}$, $j=1,2,\dots,N$, are weakly dependent mean 0 random variables. Hence, conditionally on $\lambda_{1},\dots,\lambda_{N}$, CLT applies and the sum is asymptotically Gaussian. The limit of the variance can be computed using
@@
For the remaining eigenvalues, the idea is to show that the (i+1)(i+1)st largest root of (9.8) is very close to $\lambda_{i}$ and then use the induction hypothesis. By Corollary 9.3, the (i+1)(i+1)st largest root is the unique root in the interval $(\lambda_{i+1},\lambda_{i})$ and our task is to show that it is much closer to the right end-point of this segment rather than to the left end-point.
@@
Note that for $a$ close to $\lambda_{i}$, the value of $\frac{1}{2}\bigl(a-\sqrt{a^{2}-4}\bigr)$ is close to $1/\theta_{i+1}>1/\theta_{1}$ by induction assumption (9.5). Hence, for large $N$ the left-hand side of (9.13) is negative and bounded away from zero. On the other hand $\zeta_{i}^{2}=O(\tfrac{1}{N})$. We conclude that $a$ should be smaller than $\lambda_{i}$, at distance $O(\tfrac{1}{N})$, in order for (9.13) to hold. Hence, this root $a$ satisfies (9.5).
Case 2: $\theta_{i+1}<\theta^{c}$, so that $\lambda_{i}$ is close to $\lambda_{+}$. We approximate the right-hand side of (9.8) as $N\to\infty$ near $\lambda_{+}$ by writing $\zeta_{j}^{2}\stackrel{{\scriptstyle d}}{{=}}\frac{\xi_{j}^{2}}{\sum_{\ell=1}^{N}\xi_{\ell}^{2}}$ with i.i.d. $\mathcal{N}(0,1)$ random variables $\xi_{j}$ and then using Theorem 8.20 with $\mathfrak{s}=\mathfrak{m}=1$, $h(x)=1$, whose assumptions hold by the induction hypothesis. The right-hand side of (9.8) has asymptotics $1+N^{-1/3}\cdot\mathcal{G}(b)+o(N^{-1/3})$ where the rescaled variable is $b=N^{2/3}(a-2)$. The interval $a\in(\lambda_{i+1},\lambda_{i})$ turns asymptotically into $b\in(\mathfrak{a}_{i-q+2},\mathfrak{a}_{i-q+1})$, and the equation becomes $\mathcal{G}(b)=N^{1/3}(1/\theta_{1}-1)+o(N^{-1/3})$. Since $1/\theta_{1}<1/\theta^{c}=1$, we are looking for a value of $b\in(\mathfrak{a}_{i-q+2},\mathfrak{a}_{i-q+1})$, where $\mathcal{G}(b)$ would be large and negative. Clearly, then $b$ needs to be close to the right end-point of the segment, i.e., to $\mathfrak{a}_{i-q+1}$ and we achieve (9.6).
@@
It remains to investigate $a_{q}$. Because $\mathbf{u}^{*}$ is uniformly random and $\mathbf{u}_{i}]_{i=1}^{N}$ are orthonormal, the vector $\langle\mathbf{u}_{j},\mathbf{u}^{*}\rangle^{2}$, $1\leq j\leq N$, has the same distribution as $\frac{\xi_{j}^{2}}{\sum_{\ell=1}^{N}\xi_{\ell}^{2}}$ with i.i.d. $\mathcal{N}(0,1)$ random variables $\xi_{j}$. Hence, recalling $\theta_{q}=\theta^{c}+N^{-1/3}\tilde{\theta}=1+N^{-1/3}\tilde{\theta}$, (9.1) becomes:
@@
Taylor expanding the left-hand side in small $N^{-1/3}\tilde{\theta}$ and using Corollary 8.22, we conclude that the desired root $b$ converges towards the largest root of the equation $\tilde{\theta}=\mathcal{G}(b)$. Comparing with Definition 4.2, we are done. ∎
@@ 9.2. Spiked covariance model
We explain the new parts, but are brief on the technical details repeating the previous section. At the end we also analyze $\theta_{i}\to\infty$ limits and explain how those connect to the strong signals. We start with an analogue of Proposition 9.1.
Suppose that we are given $N\times S$ matrix $\mathbf{U}$, in which the rows are indexed by $i=0,1,\dots,N-1$ and $S\geq N$. We let $\lambda_{*}$ denote the squared length of the zeroth row of $\mathbf{U}$ (treated as an SS–dimensional vector) and let $\mathbf{v}^{*}$ be the unit vector in the direction of this row. Let $\widetilde{\mathbf{U}}$ denote the $(N-1)\times S$ matrix formed by rows $i=1,2,\dots,N-1$ of $\mathbf{U}$.
We would like to connect the singular values and singular vectors of $\mathbf{U}$ to: singular values and vectors of $\widetilde{\mathbf{U}}$,  $\lambda_{*}$,  and $\mathbf{v}^{*}$. We let $(\mathbf{u}_{i},\mathbf{v}_{i},\sqrt{\lambda_{i}})$, $1\leq i\leq N-1$, be the left singular vector (of $(N-1)\times 1$ dimensions), right singular vector (of $S\times 1$ dimensions), singular value triplets for $\widetilde{\mathbf{U}}$, which means that
@@
In order to see how Proposition 9.6 is relevant for the setting of Section 2.2, we choose one index $1\leq k\leq r$ and a deterministic orthogonal matrix $O$, which maps $\mathbf{u}_{k}^{*}$ into the zeroth basis vector and orthogonal complement of $\mathbf{u}_{k}^{*}$ to the span of the basis vectors with labels $1,2,\dots,N-1$, so that
@@
where $\Omega^{\prime}$ is $(N-1)\times(N-1)$ symmetric matrix with eigenvalues $\{\theta_{i}\}_{i\neq k}$ and $1$ of multiplicity NrN-r. Conjugating with $O$ does not change the eigenvalues of the sample covariance matrix $\frac{1}{S}XX^{\mathsf{T}}$; it also does not change the fact that the columns of $X$ are i.i.d. On the other hand, after transformation by $O$, we can use Proposition 9.6 for $\mathbf{U}=\Omega X$. We reach the following:
@@
On the other hand, the value of $\lambda^{*}$ in (9.18) is the squared length of the zeroth row, $\|\mathbf{v}^{*}\|^{2}=\sum_{l=1}^{S}\theta_{k}\xi_{l}^{2}$. Finally, Proposition 9.6 deals with $\mathbf{U}\mathbf{U}^{\mathsf{T}}$, while in the corollary the matrix $XX^{\mathsf{T}}$ is divided by $S$ and, therefore, we should rescale by $S$ both $\lambda_{i}$ and $a$. Hence, dividing (9.18) by $a$ and $\theta_{k}$, we get an equivalent form of (9.19):
$\frac{1}{a}\left(\frac{\sum_{i=1}^{S}\xi_{i}^{2}}{S}+\frac{1}{S}\sum_{i=1}^{N-1}\frac{\lambda_{i}\xi_{i}^{2}}{a-\lambda_{i}}\right)=\frac{1}{\theta_{k}}.\qed$
@@
(9.21) $\lim_{N\to\infty}N^{2/3}\frac{\lambda_{i}-(1+\gamma)^{2}}{\gamma(1+\gamma)^{4/3}}$ $\stackrel{{\scriptstyle d}}{{=}}\mathfrak{a}_{i-q+1},\qquad i\geq q,$
where $\mathcal{N}(0,V(\theta_{i}))$ are independent over $i$ and with points of the Airy1 point process $\{\mathfrak{a}_{j}\}_{j\geq 1}$. In addition, (8.45) with $h(x)=1$ holds for the eigenvalues of $\frac{1}{S}XX^{\mathsf{T}}$.
@@
The final claim, (8.45) with $h(x)=1$ follows by induction on $r$ from Theorem 8.24 for $r=0$, and using Lemma 8.28 with Corollary 9.8 for the induction step. Here $\lambda_{\pm}=(1\pm\gamma)^{2}$,
(9.22) $\mu(x)\mathrm{d}x=\frac{1}{2\pi}\frac{\sqrt{(\lambda_{+}-x)(x-\lambda_{-})}}{\gamma^{2}x}\,\mathbf{1}_{[\lambda_{-},\lambda_{+}]}\,\mathrm{d}x,\qquad m(z)=\frac{z+\gamma^{2}-1-\sqrt{(z-\lambda_{+})(z-\lambda_{-})}}{2\gamma^{2}z},$
@@
We analyze the equation (9.19) for $k=q$. In this situation the asymptotics of $(\lambda_{i})_{i=1}^{N}$ is given to us by Proposition 9.9. Arguing as in the previous section, the q1q-1 largest roots of the equation are close to $\lambda_{1},\dots,\lambda_{q-1}$, resulting in (4.6). In order to establish (4.7), we need to approximate (9.19) for $a$ close to $\lambda_{+}=(1+\gamma)^{2}$ and locate the root of the equation in the $(\lambda_{q},\lambda_{q-1})$ interval. We change the variables
@@
Recognizing the constant $\kappa_{2}$ and comparing with Definition 4.2, we arrive at (4.7). ∎
We now analyse $\theta_{i}\to\infty$ limits. The following is immediate from (2.8).
Corollary 9.10.
The parameters in (4.6) for the spiked covariance model satisfy:
(9.28) $\frac{\lambda(\theta)}{\theta}=1+O\left(\frac{1}{\theta}\right),\qquad\frac{V(\theta)}{\theta^{2}}=2\gamma^{2}+O\left(\frac{1}{\theta}\right),\qquad\theta\to\infty.$
Corollary 9.10 agrees with results in the literature (e.g. Wang and Fan (2017); Cai et al. (2020); Jiang and Bai (2021); Fan et al. (2024a)) for the cases when $\theta$ grows with $N$, and shows that the approximation of Theorem 4.6 remains valid in such regimes. For instance, in the strong signal regime, $\theta_{i}=N\tilde{\theta}_{i}$ as $N\to\infty$, formally inserting $\theta_{i}=N\tilde{\theta}_{i}$ into (2.8), dividing by $\gamma N\tilde{\theta}_{i}$ and using (9.28), we get $\sqrt{S}\left(\frac{\lambda_{i}}{\theta_{i}}-1\right)\xrightarrow[]{d}\mathcal{N}(0,2),$ which precisely matches Fan et al. (2024a, Theorem 1).